# Uniform rate inflation on the brane

### v2

@@ I Introduction
$\Lambda=\frac{3M_{5}^{6}}{4\pi}.$ (8)
We recover the conventional Friedmann equation if $\rho\ll 2\Lambda$ (like our current universe). The nucleosynthesis limit implies that $\Lambda\gtrsim(1\mbox{ MeV})^{4}\sim(4.1\times 10^{-22})^{4}$. A more stringent constraint can be obtained by requiring the theory to reduce to Newtonian gravity on scales larger than $1$ mm; this gives $M_{5}\gtrsim 10^{5}\mbox{ TeV}$. which corresponds to $\Lambda\gtrsim 5.0\times 10^{-53}$ Maartens:1999hf . We consider a braneworld described by Eq. (7) in the following discussion.
@@
$V(\phi)=\frac{\Lambda c}{2}e^{\sqrt{\frac{3\lambda^{2}}{2\Lambda}}}+\frac{\Lambda}{2c}e^{-\sqrt{\frac{3\lambda^{2}}{2\Lambda}}}-\Lambda-\frac{\lambda^{2}}{2},$ (11)
where $c$ is an arbitrary (suitably rescaled) integration constant. For simplicity, we choose $c=1$ to obtain
@@
We can regard this as the low-energy limit where the brane high-energy corrections become unimportant. The result is that we recover Eq. (4) with the predictions of primordial perturbations given in Lin:2023xgs . Therefore, Eq. (12) may be regarded as the generalization of the uniform rate inflation given by Eq. (4) to the braneworld.
@@
When $\phi>\phi_{e}$, inflation is happening with $\ddot{a}>0$. We assume that somehow reheating happens after $\phi=\phi_{e}$ and recovers hot big bang. The inflaton field value $\phi_{60}$ at the number of e-folds $N=60$ (which corresponds to horizon exit or CMB scale) can be obtained by solving $\Delta\alpha=60$ that is
@@
Figure 8: The spectral index and tensor-to-scalar ratio. The constraints are from Planck:2018jri . The blue dashed (red) line corresponds to inflation ends at $\phi_{e}$ ($\phi_{0}$).
@@
In this section, we consider a perturbation of $\dot{\phi}$ of the value given by Eq. (5). Intuitively, we can see from Eqs. (3) and (7) that if somehow $\dot{\phi}$ is increased (decreased) with the same $V\phi$ and $V\phi^{\prime}$, $H$ would also increase (decrease). Therefore, the term $3H\dot{\phi}$ in Eq. (3) is increased (decreased). To compensate for this increasing (decreasing), a negative (positive) $\ddot{\phi}$ appears, and this reduces (enhances) $\dot{\phi}$ towards the value given by Eq. (5). There is a feedback mechanism.

### v3

@@ I Introduction
$\Lambda=\frac{3M_{5}^{6}}{32\pi^{2}}.$ (8)
We recover the conventional Friedmann equation if $\rho\ll 2\Lambda$ (like our current universe). The nucleosynthesis limit implies that $\Lambda\gtrsim(1\mbox{ MeV})^{4}\sim(4.1\times 10^{-22})^{4}$. A more stringent constraint can be obtained by requiring the theory to reduce to Newtonian gravity on scales larger than $1$ mm; this gives $M_{5}\gtrsim 10^{5}\mbox{ TeV}$. which corresponds to $\Lambda\gtrsim 4.2\times 10^{-64}$ Maartens:1999hf . We consider a braneworld described by Eq. (7) in the following discussion.
@@
$V=\frac{\Lambda c}{2}e^{\sqrt{\frac{3\lambda^{2}}{2\Lambda}}\phi}+\frac{\Lambda}{2c}e^{-\sqrt{\frac{3\lambda^{2}}{2\Lambda}}\phi}-\Lambda-\frac{\lambda^{2}}{2},$ (11)
where $c$ is an arbitrary (suitably rescaled) integration constant. Different choices of the constant $c$ correspond to the field redefinition, which we are free to make. Therefore, without loss of generality, we choose $c=1$ to obtain
@@
The result is that we recover Eq. (4) with the predictions of primordial perturbations given in Lin:2023xgs . Therefore, Eq. (12) may be regarded as the generalization of the uniform rate inflation given by Eq. (4) to the braneworld. From Eqs. (1) and (12), we have
$\rho=\Lambda\left(\cosh\sqrt{\frac{3\lambda^{2}}{2\Lambda}}\phi-1\right)\simeq\frac{\Lambda}{2}\left(\sqrt{\frac{3\lambda^{2}}{2\Lambda}}\phi\right)^{2}.$ (14)
Therefore, from Eq. (7), we can regard this as the low-energy limit where the brane high-energy corrections become unimportant.
@@
When $\phi>\phi_{e}$, inflation is happening with $\ddot{a}>0$. As discussed in Lin:2023xgs , after inflation, the inflaton field does not oscillate but keeps on going with a uniform rate if the potential remains the same, and the universe tends to collapse shortly. We assume that somehow reheating happens after $\phi=\phi_{e}$ and recovers hot big bang. The inflaton field value $\phi_{60}$ at the number of e-folds $N=60$ (which corresponds to horizon exit or CMB scale) can be obtained by solving $\Delta\alpha=60$ that is
@@
Figure 8: The spectral index and tensor-to-scalar ratio. The constraints are from Planck:2018jri . The blue dashed (red) line corresponds to inflation ends at $\phi_{e}$ ($\phi_{0}$). The maximum value of predicted $n_{S}$ recovers the results of Lin:2023xgs .
@@
In this section, we consider a perturbation of $\dot{\phi}$ of the value given by Eq. (5). Intuitively, we can see from Eqs. (3) and (7) that if somehow $\dot{\phi}$ is increased (decreased) with the same $V\phi$ and $V\phi^{\prime}$, $H$ would also increase (decrease). Therefore, the term $3H\dot{\phi}$ in Eq. (3) is increased (decreased). To compensate for this increasing (decreasing), a negative (positive) $\ddot{\phi}$ appears, and this reduces (enhances) $\dot{\phi}$ towards the value given by Eq. (5). There is a feedback mechanism. This is not so unexpected since it is known that slow-rolling is an attractor in phase space, and our solution is effectively a slow-roll solution.